3.5 Exercise M -Relationship between concentrations, pH an Complete the following table for solutions at 259%C acidic, basic, neutral 2 8.32 1.00 4. 5.0x 10 1.0 x 10-12 Exercise J-Identifying acidic solutions by pH Circle the most acidic pH in each group: 2.5 1.5 10 2 12 0.9 1.5 1.3 7.4 12 7 10 5. 7.5 2.4 3.2 6. 8.5 3.8 11.2 7.6
Answer #1

Answer:
Formulae and criteria applied for calculations are,

pH = -log[H+] and [H+] = 10-pH.

pOH = -log[HO-] and [HO-] = 10-pOH.

pH + pOH = 14.00

pH < 7 : Acidic

pH = 7: Neutral

pH > 7: Basic

1) Given: pH = 8.32 (pH > 7 : Basic)

\therefore pOH = 14.00 - pH = 14.00 - 8.32 = 5.68

\therefore [H+] = 10-8.32. = 4.8*10-9 M

\therefore [HO-] = 10-5.68. = 2.1*10-6 M

===================================

2) SOlution Neutral.

pH = pOH = 7

\therefore [H+] = [HO-] = 1.00*10-7 M

===========================

3) Given: pH = 1.00 (pH < 7 : Acidic)

\therefore pOH = 14.00 - pH = 14.00 - 1.00 = 13.00

\therefore [H+] = 10-1. = 1.0*10-1 M = 0.1 M

\therefore [HO-] = 10-13. = 1.0*10-13 M

===================================

4) Given: [H+] = 5.00*10-8 M

pH = -log[H+] = -log(5.00*10-8) = 7.30 (pH > 7.00: Basic)

pOH = 14.00 - 7.30 = 6.70

[HO-] = 10-pOH = 10-6.70 = 2.0*10-7 M

===============================

e) Given: [HO-] = 1.0*10-12 M

pOH = -log[HO-] = -log(1.0*10-12) = 12

pH = 14 - 12 = 2 (pH < 7: Acidic)

[H+] = 10-pH = 10-2 = 0.01 M

====================XXXXXXXXXXXXX======================